Hello,
This question is not about complex numbers but I could not find any other catagory to put it under. It's solving problems by using first-degree open sentences. The question is, "Find three consecutive even numbers whose sum is
-90"

Answer by ankor@dixie-net.com(22740)

Find three consecutive even numbers whose sum is -90
:
x + (x+2) + (x+4) = -90
3x + 6 = -90
3x = -90 - 6
3x = -96
x = -32
:
numbers would be -32, -30, -28